ID |
Date |
Author |
Type |
Category |
Subject |
6351
|
Mon Mar 5 03:50:49 2012 |
kiwamu | Update | IOO | PZT1 PITCH railing |
PZT1 started railing in the pitch direction and because of this TRY doesn't go more than 0.7. I will leave it as it is for tonight.
Tomorrow I will shift the alignment of the MC to make the PZT1 happier.
Quote from #6300 |
PZT1, the one with Koji's custom mid-HV driver (#5447), is getting degraded.
|
|
6350
|
Mon Mar 5 03:22:54 2012 |
kiwamu | Update | PSL | PMC realigned |
I realigned the steering mirrors for the PMC. The trans value went up from 0.79 to 0.83.
The misalignment was largely in the pitch direction. |
6349
|
Fri Mar 2 18:55:06 2012 |
Den | Update | PEM | seis box |
I've put the seismometer box back to the 1x1, Guralp is back under MC2. When the seismometer is not plugged in, the noise is

Now, I'm going to collect some data from GUR 1 and MC_F and see if the problem with adaptive filter (increasing errror while decreasing mu) will be gone.
|
6348
|
Fri Mar 2 18:11:50 2012 |
jamie | Summary | SUS | evaluation of eLIGO tip-tilts from LLO |
[Suresh, Jamie]
Suresh and I opened up and checked out the eLIGO tip-tilts assemblies we received from LLO. There are two, TT1 and TT2, which were used for aligning AS beam into the OMC on HAM6. The mirror assemblies hang from steel wires suspended from little, damped, vertical blade springs. The magnets are press fit into the edge of the mirror assemblies. The pointy flags magnetically attach to the magnets. BOSEMS are attached to the frame. The DCC numbers on the parts seem to all be entirely lies, but this document seems to be close to what we have, sans the vertical blade springs: T0900566
We noticed a couple of issues related to the magnets and flags. One of the magnets on each mirror assembly is chipped (see attached photos). Some of the magnets are also a bit loose in their press fits in the mirror assemblies. Some of the flags don't seat very well on the magnets. Some of the flag bases are made of some sort of crappy steel that has rusted (also see pictures). Overall some flags/magnets are too wobbly and mechanically unsound. I wouldn't want to use them without overhauling the magnets and flags on the mirror assemblies.
There are what appear to be DCC/SN numbers etched on some of the parts. They seem to correspond to what's in the document above, but they appear to be lies since I can't find any DCC documents that correspond to these numbers:
TT1: D070176-00 SN001
mirror assembly: D070183-00 SN003
TT2: D070176-00 SN002
mirror assembly: D070183-00 SN006
|
6347
|
Fri Mar 2 16:05:52 2012 |
Den | Update | SAFETY | laser safety |
Today I've attended the laser safety seminar. |
6346
|
Fri Mar 2 11:05:28 2012 |
Den | Update | PEM | seis box gain |
I've replaced R2 resistor that adjusts the gain of the AD620 amplifier. Previous value 5491Ohm, new value 464Ohm, so the gain should increase up to ~200-250. Only at the N/S 1 circuit!
LISO simulation of the circuit transfer function and noise are


LISO predicts gain ~45-46 dB = 200 and noise at the level of 10uV at 1Hz. The transfer function and noise measured are


The noise measured is 5 times higher then predicted by LISO. Though I described AD620 as an ordinary amplifier with 49.9kOhm resistor connecting output and inverted input. I specified the noise spectrum 10 nV and 1/f corner frequency 30 Hz. In the AD620 datasheet noise spectrum is 10 - 100 nV depending on the gain. However, the gain is 200 and noise spectrum should be 10 nV. May be in reality it is not the case. It also possible that the noise model used by LISO is not valid for AD620 as it is not an ordinary operational amplifier. |
6345
|
Thu Mar 1 21:48:34 2012 |
Den | Update | PEM | seis box noise |
Quote: |
The noise increased at 0.5 Hz and is pretty big. This might explain the loose of coherence at low frequencies.
|
This is because spectrum analyzer did not plot the real noise spectrum at the first few points at low frequencies. I've remeasured the noise at 1mHz - 3Hz at "output -" (TP9) and compared it to the seismometer signal

The noise seems to be much less then the signal. I've measured the noise several times and once I got a huge amount of noise

I made another measurement in some time and got the low noise again. A circuit might have a bad contact somewhere.
The plan is to change AD620 adjustable resistor (R2) from 5.49kOhm to 500Ohm to increase the gain from 20 up to 200. |
6344
|
Thu Mar 1 09:26:50 2012 |
steve | Update | SUS | SOS baffle plates are ready |
Green welding glass 7" x 9" shade #14 with 40 mm hole and mounting fixtures are ready to reduce scatter light on SOS
PEEK 450CA shims and U-shaped clips will keep these plates damped.
|
Attachment 1: 03011201.PDF
|
|
6343
|
Thu Mar 1 00:05:23 2012 |
Den | Update | PEM | seis box noise |
I've moved GUR1 seismometer from MC2 to the working tables in order not to disturb the MC while working with the seismometer box. The new place for the GUR1 for a few days is near the printer, cables and blue boxes. I've cleaned all mess and wires from the floor, so that seismometer now looks like that

I've connected 2 inputs of the N/S 1 circuit of the seismometer box with a 50 Ohm resistor and measured the noise at the output. The comparison with the seismic signal is

The noise increased at 0.5 Hz and is pretty big. This might explain the loose of coherence at low frequencies. |
6342
|
Wed Feb 29 20:27:00 2012 |
Jenne | Update | Green Locking | X green beat - found it! |
Found it!
The actual temperature of the Xend laser is 0.02 C higher than anticipated based on the formula in elog 3759. Both the PSL and the Xend laser are at their nominal diode currents (2.100 A for the PSL, 2.003 A for Xend), so the curves should be used as they are. The PSL temp (when the slow servo offset is ~0) is 31.71 C. Using curve 2 from elog3759, the Xend laser should be 37.78, which I found was +10 counts on the Xgreen slow servo offset.
Right now the Xend laser is at 37.80 C, and the beat is around 30 MHz. This is +80 counts on the Xgreen slow servo. +60 counts gave me ~80 MHz. When (a few minutes ago) the MC unlocked and relocked, it came back to a slightly different place, so the temp of the Xend laser had to go up a few 10's of counts to get the same beat freq. Right now the PSL slow servo offset is 0.076 V.
The HP8591E is set with ResBW=100kHz, Ref Level= -39dBm (so I'm not attenuating my input signal!). The largest peak I see for the beatnote is -66dBm. The nose floor around the peak is -83dBm. Trace (trace button!) A is set to MaxHoldA, and Trace B is set to ClearWriteB, so B is giving me the actual current spectrum, while A is remembering the peak value measured, so it's easier to see if I went past the peak, and just didn't see it on the analyzer.
Also, I went back and realigned the beams earlier, to ensure that there was good overlap both near the BS which combines the PSLgreen and Xgreen beams, and at the PD. The overlap I had been looking at was okay, but not stellar. Now it's way better, which made the peak easier to see. Also, also, the waveplate after the doubling oven on the PSL table is still rotated so that I get max power on the Xgreen side of things, and not much at all on the Ygreen side. I'll need to rebalance the powers, probably after we make sure we are seeing the beatnote with the BeatBox.
Next Steps:
Lay a cable from the BBPD to the BeatBox in 1X2, make the BeatBox do its thing.
Use the dichroic locking to do a sweep of the Xarm. |
6341
|
Wed Feb 29 17:32:11 2012 |
Mike | Update | Computers | PyNDS and a Plot |
Quote: |
Quote: |
Power Spectral Density plot using PyNDS, comparing 5 fast data channels for ETMX.
|
Is there any stuff to install, etc? Y'know, for those of use who don't really know how to use computers and stuff....
|
No new stuff for these computers. Everything should be installed already. |
6340
|
Wed Feb 29 04:23:14 2012 |
kiwamu | Update | LSC | REFL OSA installed |
I placed the OSA (Optical Spectrum Analyzer) on the AP table and this OSA will monitor the REFL beam.
Tomorrow I will do fine alignment of the OSA.
(some notes)
- a new 90% BS in the REFL path for limiting the REFL beam power
I installed a 90 % beam splitter in the REFL path so that this BS limits the maximum power in the downstreams because we don't want to damage any more RFPDs.
The REFL beam has a power of about 610 mW and the BS has R = 94 % (the spec says 90 +/- 4 % ), resulting in a power of ~37 mW in the transmitted light.
Then the transmitted beam goes through the combination of a half-wave plate and PBS, which allows a fine adjustment of the power.
After passing through the lambda/2 + PBS, the beam is branched to four ways and each beam goes to the REFL RFPD, i.e. REFL11, 33, 55 and 165.
In the end each RFPD receives a laser power of 9 mW at maximum, which is reasonably lower than the power rate of the photo diodes (~17 mW ).
The new OSA uses the reflected light from the 90% BS.
- Squeezed the ABSL (ABSolute length Laser) path
I squeezed the path of the ABSL in order to accommodate the OSA.
I tried to keep the same optical distances for some lenses, but I guess their mode matching must be different from what they used to be.
So be aware of it.
- Modification of the AS OSA path
I have also modified the optical path of the AS OSA because there had been an extra zig-zag path which made the path more complex in unnecessary way.
Since I have squeezed the ABSL path, it allowed me to simplify the optical path. So I modified the path.
Quote from #6336 |
I am installing an OSA on the AP table and it's ongoing.
|
|
6339
|
Wed Feb 29 01:14:40 2012 |
Suresh | Update | Electronics | REFL165 repair: Characterization |
Quote: |
The transfer function and current noise were measured. The location of the peak shifts with the amount of incident light power (RF or DC). The TF was measured at an incident 1064nm light power of 0.4 mW which produced a DC output voltage of 14 mV => DC photocurrent of 0.28 mA.
Many of the effects that Koji noted in the previous characterization are still present.
In addition I observed a shift of the peak towards lower frequencies as the RF power supplied to the AM Laser (Jenne Laser) is increased. This could create a dependance of the demodulation phase on incident RF power.
The plots are attached below.
|
[Koji, Suresh]
To determine the amount of RF power in the AM laser beam at various RF drive levels I measured the RF power out of the Newfocus 1611 PD while driving the AM laser with a Marconi. During this measurement the DC output was 2.2V. With the DC transimpedance of 10^4 and a sensitivity of 0.8 A/W we have carrier power as 0.275 mW (-5.6 dBm). [Incidentally the measured carrier power with a power meter is about 0.55 mW. Why this discrepancy?]
|
1 |
2 |
3 |
4 |
5 |
6 |
Marconi Output (dBm) |
0 |
-5 |
-10 |
-15 |
-20 |
-25 |
AG 4395 measurement (dBm) |
-8.1 |
-13.0 |
-18.0 |
-23 |
-28 |
-33 |
RF/DC ratio dB |
-2.5 |
-7.4 |
-12.4 |
-17.6 |
-22.6 |
-27.6 |
Estimation of the signal strength at the REFL165 PD:
From the 40m Sensing Matrix for DRFPMI we see that the signal strength at REFL165 in CARM is about 5x10^4 W/m. Since we expect about 0.1nm of linear range in CARM length we expect about 0.05 mW of RF power. If the (DC) carrier power is about 10 mW at the photodiode (18mW is about the max we can have since the max power dissipation is 100 mW in the diode) then the RF : DC power ratio is 5x10^-3 => -23 dB
As this is lower than the power levels at which the PD transfer function was determined and where we noted the distorsion and shift of the resonance peak, it is likely that these effects may not be seen during the normal operation of the interferometer.
The shift due to the carrier power level (DC) change may still however pose a problem through a changing demodulation phase.
|
6338
|
Wed Feb 29 01:02:06 2012 |
Den | Update | PEM | seis box measured |
I've measured the input signal to the seismic box from seismometer Guralp 1. The spectrum of the signal in the "input +" (TP 1) is

The spectrum below 1 Hz is ~250 uV/sqrt(Hz). As the input is differential, then the input voltage is 0.5 mV/sqrt(Hz). The spectrum of the "output +" signal (TP 2) is

So the gain at ~ 1Hz is ~20. I've measured the transfer function between the "input +" and "output +" (TP1 and TP2) for all 9 circuits

The channels 1-6 are of new modification and have gain ~20 at the frequencies 0.2 - 100 Hz. Below 0.2 Hz the gain is reduced. 100 Hz - cut off frequency of the low-pass filters. Meanwhile channels 7-9 (old configuration) have much more gain and 10_50 Hz filters work here.
The coherence between "input +" and "output +" (TP1 and TP2) for 9 circuits is

We can see that channel VERT 3 is very bad. For others coherence is lost below 0.2 Hz. The spectrum analyzer noise measured is ~1000 times less then the signal at these frequencies. I'll pay more attention to this loss of coherence at low frequencies. Something is noisy. |
6337
|
Wed Feb 29 00:22:35 2012 |
Suresh | Update | Electronics | REFL165 repair: Installed on the AS table |
1) The REFL165 has been replaced onto the AS table.
2) When the PD interface cable is attached the PD shows a DC out put of 6mV and does not respond to a flash light. I changed the PD interface port in the LSC rack by swapping the other end of the cable with an unused (Unidentified PD) interface cable, The PD is working fine after that. There could be a problem with some binary switch state on the PD interface where the REFL165 cable was plugged in earlier.
|
6336
|
Tue Feb 28 20:49:33 2012 |
kiwamu | Update | LSC | insalling OSA |
I am installing an OSA on the AP table and it's ongoing.
I am leaving some stuff scattered on the AP table and I will resume the work after I come back. |
6335
|
Tue Feb 28 16:44:56 2012 |
rana | Update | LSC | MICH and PRCL signals in a simulation |
Like I said, this is possible if you fail to set up the OSA to look at the sidebands at BOTH the AS and REFL ports at all times. Don't waste your time - set up an OSA permanently! |
6334
|
Tue Feb 28 16:39:25 2012 |
kiwamu | Update | LSC | MICH and PRCL signals in a simulation |
I briefly ran a Optickle code to see how the PRC macroscopic length screws up the sensing matrix in the PRMI configuration.
Especially I focused on the optimum demodulation phases for the MICH and PRCL signals to see how well they are separated in different PRC length configuration.
It seems that the demod phases for MICH and PRCL are always nicely separated by approximately 90 degree regardless of how long the PRC macroscopic length is.
If this is true, how can we have such a strange sensing matrix ??
(Simulation results)
The plots below show the simulation results. The x-axis is the macroscopic length of PRC in a range from 6.3 meter to 7.3 meter.
The y-axis is the optimum demodulation phases for MICH (blue) and PRCL (black).
The red line is the difference between the MICH and PRCL demodulation phases.
The left plot is for the REFL11 signals and the right plot is for the REFL55 signals.
When the difference is 90 degree, it means we can nicely separate the signals (i.e. REFL11I for PRCL and REFL11Q for MICH).
Obviously they are always nicely separated by ~ 90 deg.
 
Quote from #6330 |
The lock of the PRMI doesn't look healthy, especially the sensing matrix doesn't make sense at all ( #6283).
A very staring thing in the sensing matrix is that the REFL11 and REFL55 didn't show the 90 degree separation between MICH and PRCL.
|
|
6333
|
Tue Feb 28 16:31:08 2012 |
Suresh | Update | Electronics | REFL165 repair: Characterization |
The transfer function and current noise were measured. The location of the peak shifts with the amount of incident light power (RF or DC). The TF was measured at an incident 1064nm light power of 0.4 mW which produced a DC output voltage of 14 mV => DC photocurrent of 0.28 mA.
Many of the effects that Koji noted in the previous characterization are still present.
In addition I observed a shift of the peak towards lower frequencies as the RF power supplied to the AM Laser (Jenne Laser) is increased. This could create a dependance of the demodulation phase on incident RF power.
The plots are attached below. |
Attachment 1: REFL165_Characterization.pdf
|
|
Attachment 2: REFL165_response_shift.pdf
|
|
6332
|
Tue Feb 28 16:12:59 2012 |
Den | Update | Adaptive Filtering | lunch talk |
Just to be clear what I said at the meeting, I write all this down here. Adaptive filtering of real signals (MC_F and GUR1_X) with all noises inside is

This is offline filtering but with real signals and with the C-code that is compiled at the 40m now. We can reduce the MC_F signal by ~100 below 10 Hz, but the problem is that reducing the adaptation gain, the error increases. As a result when we move towards FxLMS algorithm with AA, AI and downsampling, we have to take the gain equal to ~1e-2 and we do not reduce any noise.
The second demonstration of this problem is static Wiener filtering. This is the result

We can see that adaptive filtering outperforms the "optimal" filtering. This is because an adaptive filter can follow the changes of coefficients immediately while the Wiener filter averages them. This is the mathematical formulation:
mcl_real = coeff _real* seismic_noise_real + other_noise
mcl_real - the real length of the MC,
coeff_real - real coefficients, that represent the transfer function between seismic noise and MC length,
other_noise - noise uncorrelated to the seismic noise seismic_noise_real
But in the world of our measured signals we have the equation
mcl_measured = coeff * seismic_noise_measured + other_noise
mcl_measured = TF_mcl * mcl_real
seismic_noise_measured = TF_seis * seismic_noise_real
where TF_mcl and TF_seis - transfer functions from the real world to measurements.
It seems to me that TF_mcl or TF_seis are not constants and for that reason the TF between measured seismic noise and mcl is not constant. But it is exactly what an adaptive or Wiener filter tries to define:
coeff(time) = average(coeff(time)) + delta(coeff(time))
The result of applying average(coeff) is the green line in the Figure 2 - error after applying the Wiener filtering.
delta(coeff) - the changing part of the transfer function is caught by the adaptive filtering. The lower the gain, the lower is the capability of adaptive filter to catch these changes. Theoretically. the error after applying adaptive filter can be presented like this:
E(error*error) = E(other_noises*other_noises) + 1/(2-mu)*mu*E(other_noises*other_noises) + 1/ {mu*(2-mu)} * Tr(Q) * A
where mu = adaptation gain
Q - covariance matrix of delta(coeff)
A - norm of the seismic signal
The first term in this equation is the dispersion of other noises, the second term is the error of the adaptive filter due to non-zero gain, the third term is due to the changes in the transfer function - we can see that it is proportional to 1/mu. This term explains why the error increases while mu decreases.
Now I'm looking for the part in the path of the signals where the transfer function can change. As I mentioned above, this is not a change in the real world, it is the change in the measured signlals. My first guess is the quantization error - we do not have not enough counts. If this is not the case, I'll move to other things of the signal path. |
6331
|
Tue Feb 28 15:48:32 2012 |
kiwamu | Update | LSC | installed anti-whitening filters |
I installed the rest of the precise anti-whitening filters. Now all of the LSC sensors have the right filters.
Quote from #6330 |
I found that none of the filter banks in the LSC input signals have the precise anti-whitening filters.
I installed the precise filters on REFL11, REFL33, REFL55 and AS55 based on Jenne's measurement (#4955)
|
|
6330
|
Tue Feb 28 12:00:54 2012 |
kiwamu | Update | LSC | installed anti-whitening filters |
I found that none of the filter banks in the LSC input signals have the precise anti-whitening filters.
I installed the precise filters on REFL11, REFL33, REFL55 and AS55 based on Jenne's measurement (#4955)
After installing them I briefly checked the REFL11 sensing matrix with the PRMI locked, but it didn't change so much from what I got (#6283).
But I felt that the PRMI became more robust after that ... I just felt so ...
(Background)
The lock of the PRMI doesn't look healthy, especially the sensing matrix doesn't make sense at all ( #6283).
A very staring thing in the sensing matrix is that the REFL11 and REFL55 didn't show the 90 degree separation between MICH and PRCL.
So I suspected some electronics, particularly the demodulation boards.
(What I did)
I checked the anti-whitening filters shape to see if they are ok or not.
I found that they all had the default filters of two zeros at 150 Hz and two poles at 15 Hz. So they weren't quite tuned.
I thought this could be a problem when I measure the sensing matrix because I usually excite the length DOFs at a high frequency of 283.1 Hz
and the mismatches between the anti-whitening and whitening filters may lead to something funny at such a high frequency.
So I installed the precise filters on REFL11, REFL33, REFL55 and AS55.
After that I did a orthogonality test on each I-Q pair of the demod signals to correct the D-phases and the relative gain between I and Q.
(Next ?)
Rana and I discussed the plan and decided to go back to a simple Michelson which should be easy enough to understand what is going on and should allow us a complete set of measurements.
Our big concern behind it is that we maybe locking the PRMI at a funny operation point.
In order to assess the issue I will do the following actions on the Michelson at first and then apply the same things on the PRMI later :
- Check the amount of of the sidebands using the OSA
- Check the amount of the DC light
- Check the sensing matrix to see if the absolute values in watt / meter make sense or not
- This work needs calibrations on all the demodulated board (this is equivalent to measuring the conversion losses of the mixers in the demod boards).
- Measure the contribution from the RAMs (it must be measurable by some means)
|
6329
|
Tue Feb 28 11:20:51 2012 |
steve | Update | SAFETY | safety audit 2012 |
Correction list by visiting safety committee, Haick Issaian is not shown:
1, update laser, crane operator list and post it
2, check fire extinguishers monthly, date and initials must be on the tags
3, move drinking water tower so it does not block fire extinguisher
4, post updated crane doc at cranes
5, post present phone lists at IFO room phones
6, emergency laser shutoff at the south end must be mounted without C-clamp
7, use heavy cable tie to insure position of mag-fan on cabinet top
Additional to do list:
a, safety glasses to be cleaned
b, let the electrical shop fix Rack-AC power to optical tables at the ends
c, measure transmission of laser safety glasses
d, update IFO outside door info signs
e, update laser inventory and post it
f, schedule annual crane inspection and renew maintenance contract
g, PSL enclosure inner shelf needs a good clean up so it is earthquake safe
|
Attachment 1: safety12.JPG
|
|
6328
|
Mon Feb 27 21:26:22 2012 |
Den | Update | PEM | seis box |
I did liso simulation of the circuit in the seis box. I think that AD620 (first amplifier in the circuit) noise might be much less with the signal from guralps from 0.01 Hz. Here is the TF of AD620 output / circuit input.

The noise spectrum is at this node is

The psd of the seismic noise below 1 Hz ~ 1u m/s => circuit input signal is ~1mv.
The TF of the whole circuit is

This result differs from the graph on the circuit sheet, but may be it was done before the resistor parameteres changed. Back of the envelop calculations also show that it is not possible to acheive DC gain 200 while 50-800 Hz gain = 5000. I'll check with the spectrum analyzer.
AD620 might be a weak point in the simulation since this is not a "classical" operational amplifier, it contains a resistor that adjusts the gain. During the liso simulation I assumed that we have an ordinary opamp (with noise, gain and gbw parameters taken from the real ad620 datasheet) with a resistor parallel to the opamp = 50k and a resistor before the inverted input that corrsponds to R2. In this case the gain of the simulated opamp is the same as of the real one given by the formula 1 + 49.9k / R2, though noise parameters may change. This should be also checked with the spectrum analyzer. |
6327
|
Mon Feb 27 19:04:13 2012 |
jamie | Update | CDS | spontaneous timing glitch in c1lsc IO chassis? |
For some reason there appears to have been a spontaneous timing glitch in the c1lsc IO chassis that caused all models running on c1lsc to loose timing sync with the framebuilder. All the models were reporting "0x4000" ("Timing mismatch between DAQ and FE application") in the DAQ status indicator. Looking in the front end logs and dmesg on the c1lsc front end machine I could see no obvious indication why this would have happened. The timing seemed to be hooked up fine, and the indicator lights on the various timing cards were nominal.
I restarted all the models on c1lsc, including and most importantly the c1x04 IOP, and things came back fine. Below is the restart procedure I used. Note I killed all the control models first, since the IOP can't be restarted if they're still running. I then restarted the IOP, followed by all the other control models.
controls@c1lsc ~ 0$ for m in lsc ass oaf; do /opt/rtcds/caltech/c1/scripts/killc1${m}; done
controls@c1lsc ~ 0$ /opt/rtcds/caltech/c1/scripts/startc1x04
c1x04epics C1 IOC Server started
* Stopping IOP awgtpman ... [ ok ]
controls@c1lsc ~ 0$ for m in lsc ass oaf; do /opt/rtcds/caltech/c1/scripts/startc1${m}; done
c1lscepics: no process found
ERROR: Module c1lscfe does not exist in /proc/modules
c1lscepics C1 IOC Server started
* WARNING: awgtpman_c1lsc has not yet been started.
c1assepics: no process found
ERROR: Module c1assfe does not exist in /proc/modules
c1assepics C1 IOC Server started
* WARNING: awgtpman_c1ass has not yet been started.
c1oafepics: no process found
ERROR: Module c1oaffe does not exist in /proc/modules
c1oafepics C1 IOC Server started
* WARNING: awgtpman_c1oaf has not yet been started.
controls@c1lsc ~ 0$
|
6326
|
Mon Feb 27 18:35:45 2012 |
Jenne | Update | Green Locking | X Beat Search |
Meh. I've searched in steps of 20 counts in C1:GCX-SLOW_SERVO2_OFFSET units (16 bit +\- 10V DAC, and 1GHz/V coeffecient for the Xgreen aux laser means this is ~0.6MHz per 20 count step). I went from -400cts to +800 cts and haven't found the beatnote yet. Meh.
Both PSL green and Xgreen beams are going to the Xgreen BBPD. Both beams are easily visible, so while I didn't actually measure the power, it should be sufficient. The arm is being re-locked in green for each step, but it's not locked in IR, but that doesn't matter for just finding the beatnote.
I've got the output of the BBPD directly connected to the 50 ohm input of the HP8591E spectrum analyzer, with the freq span from 10MHz to 120MHz. The BBPD is supposed to be good up to ~100MHz, so I should catch any beatnote that's there. I have to head out, so I guess I'll continue the search tomorrow.
One of Kiwamu's suggestions was that, since no one is using the Ygreen concurrent with my fiddling, I rotate the waveplate after the PSL doubling oven so that max power goes to the Xgreen path, thus giving myself a bigger signal. I'll try that tomorrow. Today, I didn't ever touch the waveplate. |
6325
|
Mon Feb 27 18:33:11 2012 |
jamie | Update | PSL | what to do with old PSL fast channels |
It appears that the old PSL fast channels never made it into the new DAQ system. We need to figure out what to do with them.
A D990155 DAQ Interface card in far right of the 1X1 PSL EuroCard ("VME") crate is supposed output various PMC/FSS/ISS fast channels, which would then connect to the 1U "lemo breakout" ADC interface chassis. Some connections are made from the DAQ interface card to the lemo breakout, but they are not used in any RTS model, so they're not being recorded anywhere.
An old elog entry from Rana listing the various PSL DAQ channels should be used as reference, to figure out which channels are coming out, and which we should be recording.
The new ALS channels will need some of these DAQ channels, so we need to figure out which ones we're going to use, and clear out the rest.
|
6324
|
Mon Feb 27 14:35:37 2012 |
Jenne | Update | Green Locking | PSL Beat Setup |
Xarm is aligned for both IR and green.
Here is a photo of the beam paths of the PSL beat setup. I want to make sure that the X-green BBPD sees a nice beam from both the PSL and the Xarm, without disturbing the currently working Y setup. I keep getting confused with all the beamsplitters, especially the green PBSes, which operate at ~56deg, not 45deg, so I made a diagram.

|
6323
|
Mon Feb 27 14:35:22 2012 |
steve | Update | SUS | oplev YAW transfer functions |
The BS and the PRM have 3.3 Hz resonant gain filters that kill the phase margins.
|
Attachment 1: oplYAWtransf.pdf
|
|
6322
|
Mon Feb 27 10:21:37 2012 |
steve | Update | SAFETY | safety audit tomorrow morning |
Quote: |
Emergency exit lights were inspected: 2 out of 13 batteries have to be replaced
One of the Halon fire extinguishers needs to be recharged out of 8
Please do participate in preparation for the upcoming safety audit on Feb 28
|
Batteries replaced and cylinder recharged. Please clean up your experimental set up if it is blocking breakers or entry way etc.
I will start the final clean up 2pm today.
|
6321
|
Sat Feb 25 14:27:26 2012 |
kiwamu | Update | LSC | glitches in the RFPD outputs |
Last night I took a closer look at the LSC analog signals to find which components are making the glitches.
I monitored the RFPD output signals and the demodulated signals at the same time with an oscilloscope when the PRMI was kept locked.
Indeed the RFPD outputs have some corresponding fast signals although I only looked at the RELL11 I and Q signals.
(REFL33 didn't have sufficiently a high SNR to see the glitches with the oscilloscope.)
I will check the rest of channels. |
6320
|
Sat Feb 25 00:37:42 2012 |
kiwamu | Update | SUS | oplev spectra during PRMI lock |
Somehow the angular stability of the central part have not been so great.
Also the angular motions look fluctuating a lot and they seem to be related with the glitches.
I took the oplev spectra when the PRMI is locked and unlocked to see whether if something obviously crazy is going on or not.
They seem ok to me except that the PRM pitch shows an extra bump at around 2-3 Hz when the PRMI is locked. But I don't think it's prominent.
- The attached files show the oplev spectra. When the PRMI is locked the PRM and both ITMs are under the length control.
(red) pitch when PRMI is locked
(blue) yaw when PRMI is locked
(orange) pitch without any length controls
(cyan) pitch without any length controls
|
Attachment 1: oplev_PRMI.pdf
|
|
6319
|
Fri Feb 24 23:14:09 2012 |
kiwamu | Update | CDS | tdsavg went crazy |
I found that the LSCoffset script didn't work today. The script is supposed to null the electrical offsets in all the LSC channels.
I went through the sentences in the script and eventually found that the tdsavg command returns 0 every time.
I thought this was related to the test points, so I ran the following commands to flush all the test point running and the issue was solved.
[term]> diag
[diag]>open
[diag]> diag tp clear *
EDIT, JCD 11June2012: 3rd line there should just be [diag]> tp clear * |
6318
|
Fri Feb 24 19:25:43 2012 |
jamie | Update | LSC | ALS X-arm beatbox added, DAQ channels wiring normalized |
I have hooked the ALS beatbox into the c1ioo DAQ. In the process, I did some rewiring so that the channel mapping corresponds to what is in the c1gcv model.
The Y-arm beat PD is going through the old proto-DFD setup. The non-existant X-arm beat PD will use the beatbox alpha.
Y coarse I (proto-DFD) --> c1ioo ADC1 14 --> C1:ALS_BEATY_COARSE_I
Y fine I (proto-DFD) --> c1ioo ADC1 15 --> C1:ALS_BEATY_FINE_I
X coarse I (bbox alpha)--> c1ioo ADC1 02 --> C1:ALS_BEATX_COARSE_I
X fine I (bbox alpha)--> c1ioo ADC1 03 --> C1:ALS_BEATX_FINE_I
This remapping required coping some filters into the BEATY_{COARSE,FINE} filter bank. I think I got it all copied over correctly, but I might have messed something up. BE AWARE.
We still need to run a proper cable from the X-arm beat PD to the beatbox.
I still need to do a full noise/response characterization of the beatbox (hopefully this weekend). |
6317
|
Fri Feb 24 19:18:28 2012 |
kiwamu | Update | LSC | Y arm + PRMI : how they should look like |
I calculated how the DC signals should look like in the Y arm PRMI configuration.
The expected signals are overlaid in the same plot as that of shown in #6313.
You can see there are disagreements between the observed and expected signals in the plot below at around the time when the arm is brought to the resonance.
(expected behaviors)
- TRY: At the end it should be at 1 (remember TRY is normarlized) and should not go more than that, since the power-recycling is in a weird situation and it is not fully recycling the power.
- ASDC: It should become brighter at the end because the arm cavity flips the sign of the reflected light and hence the dark port must be on a bright fringe.
- REFLDC: It will decrease a little bit because the arm cavity and MICH try to suck some amount of the power into the interferometer.

Quote from #6313 |
The figure below shows the time series of the Y arm + PRMI trail.
|
|
6316
|
Fri Feb 24 18:59:04 2012 |
Jenne | Update | Computers | PyNDS and a Plot |
Quote: |
Power Spectral Density plot using PyNDS, comparing 5 fast data channels for ETMX.
|
Is there any stuff to install, etc? Y'know, for those of use who don't really know how to use computers and stuff.... |
6315
|
Fri Feb 24 18:37:13 2012 |
rana | Update | LSC | Y arm + PRMI part II |
Quote: |
I tried the Yarm + PRMI configuration again.
The PRMI part was locked, but it didn't stay locked during the Y arm was brought to the resonance point.
|
Isn't the point that the 11 and 55 MHz signals have the carrier effect, but the 3f signals are better? |
6314
|
Fri Feb 24 16:10:48 2012 |
mike | Update | Computers | PyNDS and a Plot |
Power Spectral Density plot using PyNDS, comparing 5 fast data channels for ETMX.
**EDIT** Script here:
import nds
import numpy as np
import matplotlib.pyplot as plt
import time
daq=nds.daq('fb', 8088)
channels=daq.recv_channel_list()
e=0
start=int(time.time()-315964819)
rqst=['C1:SUS-ETMX_SENSOR_UR','C1:SUS-ETMX_SENSOR_UL','C1:SUS-ETMX_SENSOR_LL','C1:SUS-ETMX_SENSOR_LR','C1:SUS-ETMX_SENSOR_SIDE'] #Requested Channels
for c in channels:
if c.name in rqst:
daq=nds.daq('fb', 8088)
data=daq.fetch(start-100, start, c.name)
vars()['psddata'+str(e)], vars()['psdfreq'+str(e)]=plt.psd(data[0],NFFT=16384,Fs=c.rate)
vars()['label'+str(e)]=c.name
e+=1
plt.figure(1)
plt.clf()
plt.title('PSD Comparison')
plt.grid(True, which='majorminor')
plt.xlabel(r'Frequency $Hz$')
plt.ylabel(r'Decibels $\frac{dB}{Hz}$')
for x in np.arange(0,e):
plt.loglog(psdfreq0, 10*vars()['psddata'+str(x)], label=vars()['label'+str(x)])
plt.legend()
plt.show() |
Attachment 1: PSD_Comparison.png
|
|
6313
|
Fri Feb 24 15:01:31 2012 |
kiwamu | Update | LSC | Y arm + PRMI part II |
The figure below shows the time series of the Y arm + PRMI trail.

(Top plot )
Normalized TRY (intracavity power). It is normalized such that it shows 1 when the arm is locked with the recycling mirrors misaligned.
(Middle plot)
ASDC and REFLDC in arbitrary unit.
(Bottom plot)
The amount of the arm length detuning observed at the fine frequency discriminator.
(Sequence)
At t = 20 sec, the amount of detuning was adjusted so that the cavity power goes to the maximum. At this point the PRM was misaligned.
At t = 30 sec, the cavity length started being slowly detuned to 10 nm. As it is being detuned the intracavity power goes down to almost zero.
At t = 45 sec, the alignment of PRM was restored. Because of that, the REFLDC and ASDC diodes started receiving a large amount of light.
At t = 85 sec, the PRCL and MICH were locked. The REFLDC signal became a high value as the carrier light is mostly reflected. The ASDC goes to a low value as the MICH is kept in the dark condition.
At t = 100 sec, the length started being slowly back to the resonance while the PRMI lock was maintained.
At t = 150 sec, the lock of the PRCL and MICH were destroyed. With the arm fully resonance, I wasn't able to recover the PRMI lock with the same demod signals.
Quote from #6310 |
I tried the Yarm + PRMI configuration again.
The PRMI part was locked, but it didn't stay locked during the Y arm was brought to the resonance point.
I will post the time series data later.
|
|
6312
|
Fri Feb 24 08:06:52 2012 |
steve | Update | SUS | sus restored |
Quote: |
The following optics were kicked:
MC1 MC2 MC3 ETMX ETMY ITMX ITMY PRM SRM BS
Fri Feb 24 04:11:15 PST 2012
1014120690
Steve (or anyone), can you restore the watchdogs when you come to the lab in the morning ?
|
Kiwamu (or whoever is here last tonight): please run the free-swing/kick script (/opt/rtcds/caltech/c1/scripts/SUS/freeswing) before you leave, and I'll check the matrices and update the suspensions tomorrow morning.
|
|
All suspentions were restored and MC locked. PRM side osem RMS motion was high.
Atm2, Why the PRM is 2x as noisy as the SRM ? |
Attachment 1: freePRM.png
|
|
Attachment 2: noisyPRM.png
|
|
6311
|
Fri Feb 24 04:12:44 2012 |
kiwamu | Update | SUS | freeswing test |
The following optics were kicked:
MC1 MC2 MC3 ETMX ETMY ITMX ITMY PRM SRM BS
Fri Feb 24 04:11:15 PST 2012
1014120690
Steve (or anyone), can you restore the watchdogs when you come to the lab in the morning ?
|
Kiwamu (or whoever is here last tonight): please run the free-swing/kick script (/opt/rtcds/caltech/c1/scripts/SUS/freeswing) before you leave, and I'll check the matrices and update the suspensions tomorrow morning.
|
|
6310
|
Fri Feb 24 03:58:13 2012 |
kiwamu | Update | LSC | Y arm + PRMI part II |
I tried the Yarm + PRMI configuration again.
The PRMI part was locked, but it didn't stay locked during the Y arm was brought to the resonance point.
I will post the time series data later.
(locking of the PRMI part)
Tonight I was able lock the PRMI when the arm was off from the resonance by 10 nm (#6306).
This time I used REFL11Q to lock the MICH instead of the usual AS55Q because the MICH didn't stay locked with AS55Q for some reason.
The PRCL was held by REFL33I as usual.
Also I disabled the power normalization for the error signals because it could do something bad during the Y arm is borough to the resonance.
In order to reduce the number of the glitches, PRM was slightly misaligned because I knew that the lower finesse gives fewer glitches. |
6309
|
Thu Feb 23 10:53:54 2012 |
steve | Update | SUS | oplev PIT transfer functions |
Oplev transfer functions PIT UGF were optimized to be at 2-3 Hz with 60 degree minimum phase margin by adjuting oplev gains.
Additional Notes by KI:
- The PRM oplev has a tailored 3.3 Hz resonant gain in order to calm down a wobble during the lock acquisitions.
- Also in the PRM oplev a 35 Hz elliptic cut-off filter wasn't activated at the time when Steve measured it.
- In both ITMs, elliptic cut-off filters seem to have higher corner frequencies compered with the others.
- I guess these settings are from the old days.
- ETMs and ITMs have whitening filters while the rest of the suspensions don't.
- Without the whitening filters, normally the signals above 30 Hz are covered by some electrical noises or perhaps He-Ne laser intensity noise (#5630).
- This is why we usually use the 35 Hz elliptic filters to roll off the control noises.
- Since the ETMs and ITMs have whitening filters they potentially can have slightly higher corner frequencies in the elliptic filters.
- Of course the corner frequencies need to be re-designed in terms of the amount of noise injection to the longitudinal motion.
|
Attachment 1: oplPITtrans.pdf
|
|
6308
|
Thu Feb 23 09:09:33 2012 |
steve | Update | SAFETY | safety checks |
Emergency exit lights were inspected: 2 out of 13 batteries have to be replaced
One of the Halon fire extinguishers needs to be recharged out of 8
Please do participate in preparation for the upcoming safety audit on Feb 28 |
6307
|
Thu Feb 23 02:20:07 2012 |
Zach | Update | SUS | wacky state of SUS input matrices |
This reminds me that the whole Dr. SUS situation never got taken care of. Where I left off, I was having issues pulling 40m data with NDS2 (which is what all the diagonalization scripts use).
What is the deal with 40m+NDS2? If it is till no-go, can we have a consensus on whether this is too important to wait for? If so, I will rewrite the scripts to use NDS and we can upgrade to NDS2 once we can prove we know how to use it.
Quote: |
While Kiwamu and I were trying to investigate the the vertex glitches we were noticing excess noise in ITMX, which Kiwamu blamed on some sort of bad diagonalization. Sure enough, the ITMX input matrix is in the default state [0], not a properly diagonalized state. Looking through the rest of the suspensions, I found PRM also in the default state, not diagonalized.
We should do another round of suspension diagonalization.
Kiwamu (or whoever is here last tonight): please run the free-swing/kick script (/opt/rtcds/caltech/c1/scripts/SUS/freeswing) before you leave, and I'll check the matrices and update the suspensions tomorrow morning.
[0]
0.25 |
0.25 |
0.25 |
0.25 |
0 |
1.66 |
1.66 |
-1.66 |
1.66 |
0 |
1.66 |
-1.66 |
-1.66 |
1.66 |
0 |
0 |
0 |
0 |
0 |
1 |
|
|
6306
|
Wed Feb 22 19:45:33 2012 |
kiwamu | Update | LSC | how much length offset do we need ? |
I did a quick calculation to see if the offset of the arm length which I tried last night was reasonable or not.
The conclusion is that the 20 nm offset that i tried could be a bit too close to a resonance of the 55 MHz sidebands.
A reasonable offset can be more like 10 nm or so where the phases of all the laser fields don't get extra phases of more than ~ 5 deg.
The attached plot shows where the resonances are for each sideband as a function of the displacement from the carrier's resonance.

The red solid line represent the carrier, the other solid lines are for the upper sidebands and the dashed lines are for the lower sidebands.
The top plot shows the cavity power and the bottom plot shows how much phase shift the fields get by being reflected by the arm cavity.
Apparently the closest resonances to the the main carrier one are that of the 55 MHz sidebands, and they are at +/- 22 nm.
So if we displace the arm length by 22 nm, either of the 55 MHz sidebands will enter in the arm cavity and screw up the sensing matrix for the 55 MHz family.
Quote from #6304 |
In all the configurations I displaced the Y arm by 20 nm from the resonance.
|
|
6305
|
Wed Feb 22 16:55:16 2012 |
Jamie | Update | SUS | wacky state of SUS input matrices |
While Kiwamu and I were trying to investigate the the vertex glitches we were noticing excess noise in ITMX, which Kiwamu blamed on some sort of bad diagonalization. Sure enough, the ITMX input matrix is in the default state [0], not a properly diagonalized state. Looking through the rest of the suspensions, I found PRM also in the default state, not diagonalized.
We should do another round of suspension diagonalization.
Kiwamu (or whoever is here last tonight): please run the free-swing/kick script (/opt/rtcds/caltech/c1/scripts/SUS/freeswing) before you leave, and I'll check the matrices and update the suspensions tomorrow morning.
[0]
0.25 |
0.25 |
0.25 |
0.25 |
0 |
1.66 |
1.66 |
-1.66 |
1.66 |
0 |
1.66 |
-1.66 |
-1.66 |
1.66 |
0 |
0 |
0 |
0 |
0 |
1 |
|
6304
|
Wed Feb 22 13:28:22 2012 |
kiwamu | Update | LSC | Y arm + central part locking |
Last night I tried the "Y arm + central part" locking again. Three different configuration were investigated :
- Y arm + DRMI
- Y arm + PRMI
- Y arm + MICH
In all the configurations I displaced the Y arm by 20 nm from the resonance.
As for the DRMI and PRMI configurations I wasn't able to acquire the locks.
As for the MICH configuration, the MICH could be locked with AS55. But after bringing the Y arm to the resonance point the lock of MICH was destroyed. |
6303
|
Wed Feb 22 01:53:57 2012 |
kiwamu | Update | LSC | update on glitch table |
I tried SRMI. The glitch rate wasn't as high as that of PRMI but it happened once per 10 sec or so.
|
Yarm
(POY11 -->
ETMY)
|
Xarm
(POX11 --> ETMX)
|
MICH
(AS55-->BS)
or
(AS55 --> ITMs)
|
Half PRMI
(REFL11 --> PRM)
or
(REFL33 --> PRM)
|
low finesse PRMI
(ASDC --> ITMs)
(REFL33 --> PRM)
|
PRMI (carrier)
(AS55 --> ITMs)
(REFL33 --> PRM)
|
PRMI (sideband)
(AS55 --> ITMs)
(REFL33 --> PRM)
|
SRMI(NEW)
(AS55-->ITMs)
(REFL11I --> SRM)
|
DRMI
|
AS55 |
NO |
NO |
NO |
NO |
glitch (depends on finesse)
|
glitch |
glitch |
glitch |
glitch |
REFL11 |
NO |
NO |
NO |
NO |
glitch (depends on finesse)
|
glitch |
glitch |
glitch |
glitch |
REFL33 |
NO |
NO |
NO |
NO |
- |
glitch |
glitch |
glitch |
glitch |
REFL55 |
NO |
NO |
NO |
NO |
glitch(depends on finesse) |
glitch |
glitch |
glitch |
glitch |
REFL165 |
NO |
NO |
NO |
- |
- |
- |
- |
- |
- |
POX11 |
- |
NO |
NO |
NO |
- |
glitch |
glitch |
- |
glitch |
POY11 |
NO |
- |
NO |
NO |
- |
glitch |
glitch |
- |
glitch |
POP55 |
- |
- |
- |
- |
- |
- |
- |
|
- |
|
|
|
|
|
|
|
|
|
|
Quote from #6284 |
I updated the table which I posted some time ago (#6231). The latest table is shown below.
It seems that the glitches show up only when multiple DOFs are locked.
|
|
6302
|
Tue Feb 21 22:06:18 2012 |
jamie | Update | LSC | beatbox DFD installed in 1X2 rack |
I have installed a proto version of the ALS beatbox delay-line frequency discriminator (DFD, formally known as MFD), in the 1X2 rack in the empty space above the RF generation box.
That empty space above the RF generation box had been intentionally left empty to provide needed ventilation airflow for the RF box, since it tends to get pretty hot. I left 1U of space between the RF box and the beatbox, and so far the situation seems ok, ie. the RF box is not cooking the beatbox. This is only a temporary arrangement, though, and we should be able to clean up the rack considerably once the beatbox is fully working.
For power I connected the beatbox to the two unused +/- 18 V Sorensen supplies in the OMC power rack next to the SP table. I disconnected the OMC cable that was connected to those supplies originally. Again, this is probably just temporary.
Right now the beatbox isn't fully functioning, but it should be enough to use for lock acquisition studies. The beatbox is intended to have two multi-channel DFDs, one for each arm, each with coarse and fine outputs. What's installed only has one DFD, but with both coarse and fine outputs. It is also intended to have differential DAQ outputs for the mixer IF outputs, which are not installed in this version.
The intended design was also supposed to use a comparator in the initial amplification stages before the delay outputs. The comparator was removed, though, since it was too slow and was limiting the bandwidth in the coarse channel. I'll post an updated schematic tomorrow.
I made some initial noise measurements: with a 21 MHz input, which corrseponds to a zero crossing for a minimal delay, the I output is at ~200 nVrms/\sqrt{Hz} at 5 Hz, falling down to ~30 nVrms about 100 Hz, after which it's mostly flat. I'll make calibrated plots for all channels tomorrow.
The actual needed delay lines are installed/hooked up either. Either Kiwamu will hook something up tonight, or I'll do it tomorrow. |